At the city museum, child admission is $5.20 and adult admission is $8.50. On Monday, 180 tickets were sold for a total sales of $1216.50. How many child tickets were sold that day?
Answer by rfer(16322)

a+c=180
c=180-a
5.20c+8.50a=1216.50
5.20(180-a)+8.50a=1216.50
936-5.20a+8.50a=1216.50
3.30a=1216.50-936
a=280.50/3.30
a=85 adults
c=180-85
c=95 children